OpenStudy (anonymous):

Write an equation of the line that passes through (1,5) and (-1, 3). y = x + 9 y = x + 4 y = -x + 4 y = -x + 9

OpenStudy (welshfella):

one way to do this is to plug in values of x and y and see which fit

OpenStudy (welshfella):

for example Option 1 try x = 1 , y = 5:- y = x + 9 5 = 1 + 9 - does not fit so Option 1 is not correct

OpenStudy (welshfella):

Note both points (1,5) and (-1,3) must fit

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so its b

OpenStudy (welshfella):

yes its b